# and quantum information science

### v1

@@ 7 Quantum complexity and space-time: a more concerted approach
In 2009, several connections between geometry and entanglement were established, including a schematic relation of entanglement in AdS spacetime to tensor networks [4]. Hartman and Maldacena computed the entanglement entropy’s dynamics in excited states of a CFT on an infinite space, using black-hole interiors [204, 205]. Specifically, they showed how the interior’s growth with time, as quantified by the area of the Hubeny-Rangamani-Takayanagi (HRT) surface, predicts a linear increase of the entanglement entropy, followed by a saturation, once the black hole reaches effective equilibrium [204]. The interior keeps growing long after the entanglement saturates. Therefore, Susskind suggested that the growth of the black hole’s interior should be encoded via the AdS/CFT correspondence by a dual quantum mechanical quantity that increases for a time exponential in the system’s size. The candidate dual quantity was quantum complexity [206].
To make the relation more precise, it was important to identify geometrical quantities that captured the interior’s linear growth and switchback effect exhibited by complexity in quantum circuits (see Sec. 4). The first proposal [^22] The term holographic conjectures appears interchangeably with holographic proposals in many texts on the subject, to emphasize that the duality between holographic complexity and complexity of quantum systems is in general conjectural (except in special cases, see below). was the wormhole’s volume, suggested by the ER=EPR principle [205]. This idea was later refined, leading to several alternative conjectures that we present in Sec. 7.1. All these holographic proposals reproduce linear growth and the switchback effect [19, 20, 21, 22, 23, 24], as we show in Sec. 7.2. There, we also discuss a two-dimensional case where holographic complexity exhibits an expected saturation at exponential time in the black hole entropy. This saturation of complexity was discussed in the context of circuit dynamics in Sec. 4.1. In addition to the time-dependent properties of quantum complexity, we are also interested in the structure of UV divergences of the geometrical quantities. The time-independent properties of the holographic complexity conjectures are analyzed in Sec. 7.3. By now, holographic complexity has been investigated in a plethora of cases, including subregions, geometries with defects, black holes in de Sitter space, and more. We summarize the main results obtained from these investigations in Sec. 7.4.
Before proceeding, let us stress that the gravitational quantities associated with the holographic complexity proposals are all well-defined geometrical measures. However, this line of research suffered for a while from a major issue: there was no precise quantitative matching between a gravitational proposal and a dual field-theory version of complexity. Most computations involved a qualitative comparison with quantum circuit counting arguments, tensor-networks and free-field-theory results. Section 7.5 collects quantitative matches between the boundary complexity and the bulk observables. These include: the relation between the size of an ERB in two-dimensional JT gravity and a double-scaled SYK model’s spread complexity (see Sec. 2 for its definition) [26, 27], a relation between Krylov complexity and the momentum of a particle in AdS spacetime, and results regarding Nielsen’s complexity in CFT. Nonetheless, matching geometrical observables in the gravitational setting with quantum complexity in field theory is difficult. This difficulty led to questions about the foundations of quantum computation in the context of the AdS/CFT correspondence. Pseudorandomness, addressed in Sec. 7.6, can be used to argue that entries in the holographic dictionary can be hard to compute, or alternatively, that quantum gravity solves problems efficiently that take a quantum computer exponentially long times. This unexpected line of research was revealed through the study of quantum complexity.
@@ 7.1 Holographic complexity proposals
We begin by defining the holographic complexity proposals, depicted in Fig. 10. We describe their time evolutions, including linear growth and the switchback effect, in Sec. 7.2.
@@
The complexity equals volume (CV) proposal [19, 18] declares that holographic complexity is proportional to the maximal volume $\mathcal{V}$ of a codimension-one spacelike surface $\mathcal{B}$ anchored at the $(d-1)$-dimensional boundary slice $\Sigma_{\rm CFT}$ where a CFT state is defined (see fig. 10(a)), i.e.,
@@
where $\ell_{\rm bulk}$ is a length scale introduced to maintain dimensional consistency. In the case of AdS spacetime, $\ell_{\rm bulk}$ is often identified with the AdS radius $L$. The maximal surface anchored on both boundaries of an eternal black hole penetrates the black hole interior, but avoids the singularity and regions with high curvature. This makes the maximal volume’s value reliable in the semiclassical approximation, without the need to make further assumptions.
@@
Each of the previous proposals presents ambiguities. CV and CV2.0 require a length scale $\ell_{\rm bulk}$ in the holographic dictionary. The CA proposal contains contributions from the null boundaries of the WDW patch, whose normal vectors depend on an arbitrary normalization constant. The last dependence can be removed by including certain counterterms in the gravitational action, but those in turn come equipped with their own (arbitrary) length scale, which we denote $L_{\rm ct}$ [207]. When studying black holes, the CA and CV2.0 proposals probe regions near the black hole singularity. A delicate cancellation between the diverging blackening factor and the vanishing volume of the sphere in the angular directions near the singularity leads to a result which does not diverge. Nevertheless, its validity in the semiclassical approximation is uncertain since the region near the singularity is not well understood in this limit.
The above ambiguities were taken as features of the holographic complexity proposal, rather than bugs, since computational complexity is also affected by several ambiguities: for instance, the choice of allowed gates in a circuit, the tolerance to reach a target state, the cost of gates etc., see Sec. 4. This perspective was taken a step further with the recent proposal that complexity equals anything (CAny), i.e., complexity may be dual to any one out of an infinite class of geometric observables. CAny requires a codimension-zero bulk region $\mathcal{M}$ with future (past) boundary $\Sigma_{+}(\Sigma_{-})$ anchored at the CFT slice $\partial\Sigma_{\pm}=\Sigma_{\rm CFT}$. One defines the following functional (see fig. 10(c)) [23, 24, 35]
@@
where the shape of the boundaries $\Sigma_{\pm}$ is varied. When more than one solution exist, the one with maximal $W_{G_{2},F_{2,\pm}}$ is chosen. Once the unique codimension-zero region (¯,Σ¯±)¯subscript¯$(\bar{\mathcal{M}},\bar{\Sigma}_{\pm})$ solving Eq. (105) is identified, holographic complexity is computed by the observable
$\mathcal{C}_{\rm Any}[G,F_{1,\pm},\bar{\mathcal{M}}](\Sigma_{\rm CFT})=\frac{1}{G_{N}\ell_{\rm bulk}^{2}}\int_{\bar{\mathcal{M}}}d^{d+1}x\,\sqrt{-g}\,G_{1}(g_{\mu\nu})$ (106)
@@
where $G_{1},F_{1,\pm}$ are scalar functions. In summary, the CAny conjecture consists of two steps: an extremization governed by the function (104) to select the relevant bulk region in the geometry, followed by the computation of the gravitational observable in Eq. (106). CAny is a large class of physical observables, compatible with qualitative features of complexity (such as a linear growth and the switchback effect), that can be built out of the building blocks available in the gravitational system.
@@
CV and CV2.0 are recovered by making different terms in Eq. (107) dominant. The CV prescription is recovered by keeping $\alpha_{-},\alpha_{\rm B}$ constant while performing the limit $\mathcal{C}_{V}=\lim_{\alpha_{+}\rightarrow\infty}\left(\alpha_{+}^{-1}\mathcal{C}_{\rm gen}\right)\,.$ In other words, only the future codimension-one surface $\Sigma_{+}$ survives, the extremization problem imposes that the slice has maximal volume, and the physical observable is the volume itself. To recover CV2.0, one fixes the bulk coefficient $\alpha_{\rm B}$ and computes $\mathcal{C}_{2.0V}=\lim_{\alpha_{\pm}\rightarrow 0}\left(\alpha_{B}^{-1}\mathcal{C}_{\rm gen}\right)\,.$ In this limit, the boundary surfaces $\Sigma_{\pm}$ are pushed towards the null surfaces originating from the CFT slice $\Sigma_{\rm CFT}$, thus composing the WDW patch. One can also recover CA from CAny, but now the functionals (104) and (106) should differ from each other. Specifically, one needs to extremize $\mathcal{C}_{\rm gen}$ in Eq. (107) in the limit $\alpha_{\pm}\rightarrow 0$ to identify the WDW patch, but use the gravitational action (102) as the observable which computes complexity.
@@ 7.2 Time-dependent properties of the holographic proposals
In this subsection, we show that all the holographic complexity proposals defined in Sec. 7.1 exhibit the advocated universal behaviour, i.e., linear growth at late times and the switchback effect in a black-hole background.
@@
where $r_{h}$ is the horizon radius. This two-sided geometry is dual to the thermofield-double state $|\mathrm{TFD}\rangle$ of two decoupled boundary CFTs (see Eq. (5)). The isometries of the background (108) allow to express the complexity observable $\mathcal{C}_{\rm Any}$ in Eq. (106) as follows,
@@
where $\cdot\coloneqq d/d\sigma$ ($\sigma$ is an intrinsic radial parameter), $V_{x}=\int d^{d-1}x$ is the transverse volume, and $v=t+r^{*}(r)$ is an infalling null coordinate defined in terms of $dr^{*}=dr/f(r)$.
Since $v$ is cyclic for $\mathcal{L}_{\varepsilon}$ in Eq. (109), [^23] This means that $\mathcal{L}_{\varepsilon}$ depends on $\dot{v}$ but not on $v$. the conjugate momenta $P_{v}^{\varepsilon}=\partial\mathcal{L}_{\varepsilon}/\partial\dot{v}$ are conserved along the profiles of $\Sigma_{\varepsilon}$. The late time regime $t\rightarrow\infty$ is achieved when each of the codimension-one surfaces $\Sigma_{\pm}$ hug a final slice at constant radial coordinate, such that the momenta approach constant values, up to exponentially-suppressed terms. This leads to the linear growth [23, 24]
@@
where $W_{L}$ are low-energy perturbations acting on the left boundary. These perturbations carry energies of the order of the black hole’s temperature (thermal-scale perturbations), which is assumed to be much smaller than the black hole’s mass. Let us assume that $|t_{k+1}-t_{k}|>t_{*}$, where $t_{*}$ is the scrambling time, and the label $k=0,1,\dots,n+1$ in the inequalities includes the boundary times via the definitions $t_{0}=-t_{R}$ and $t_{n+1}=t_{L}$. We also assume that, in the sequence of times $t_{0},\ldots t_{n+1}$ there are $n_{\rm sb}$ switchbacks (also called time folds): in other words, the absolute values $|t_{k+1}-t_{k}|$ change sign $n_{\rm sb}\leq n$ times. To keep the bulk solutions as simple as possible, we assume that the perturbations are approximately spherically symmetric. The bulk geometry dual to the perturbed state (111) can be modeled by a black hole with $n\geq 1$ shock waves produced by null matter. In this setting, any CAny observable in the late time limit reads [24]
@@
The contribution $-2n_{\rm sb}t_{*}$ reflects the spreading of the perturbations into the system during a time scale measured by the scrambling time $t_{*}=\frac{1}{2\pi T}\log\left(M/E\right)\approx\frac{1}{2\pi T}\log S$, where $E\ll M$ is the energy of the shock, and $T$ the black hole’s temperature. This characterizes the switchback effect for the CAny observables. At late boundary times $(t_{L},t_{R})$, holographic complexity will grow linearly with the sum of the boundary times $t_{L}+t_{R}$, but with a delay induced by the scrambling time of the perturbation. This phenomenon was already observed in earlier studies of the CV, CV2.0 and CA proposals [19, 21, 22]. This should not come as a surprise. As we said, those proposals can be obtained from CAny by appropriate limiting procedures.
@@
Let us discuss other refined properties of the holographic conjectures, associated with early and intermediate times compared to the scrambling time and the inverse temperature scale. It was originally speculated that the black hole mass $M$ should provide an upper bound on the complexity rate $d\mathcal{C}/dt\leq 2M/\pi$, based on Lloyd’s bound on the computational speed [212]. This statement was supported by the the full time evolution of CV in an eternal AdS black hole background, since the volume rate showed a monotonically increasing rate approaching a constant value from below at late times (e.g., see Fig. 7 in [213]). This constant exactly saturates Lloyd’s bound for large black holes. However, the rate of CA reached a maximum at finite time before approaching from above a constant final value saturating the Lloyd’s bound (e.g., see Fig. 22 in [213]). Therefore, the intermediate regime in the evolution of CA violated the bound. Since then, many other counterexamples were found [214, 215, 216, 217, 218, 219, 220, 221, 222, 223]. It was argued in [224] that a black hole should be modeled by simple gates (i.e., close to the identity), thus being incompatible with the hypotheses used to derive the Lloyd’s bound. The adaptation of Lloyd’s bound to holographic complexity is an open problem.
@@
Next, one can study the reaction of the holographic observables to shock waves corresponding to the insertion of null matter at arbitrary boundary time [209, 43]. Earlier, we assumed that the insertion times $t_{k}$ of the perturbations (which included the boundary times $t_{L},t_{R}$) were separated by an interval larger than $t_{*}$, see the discussion below equation (111). Here we will not make this assumption. For simplicity, let us fix $t_{L}=t_{R}=0$. We consider the case of a single time fold at the (negative) time $-t_{w}$ when the shock is inserted, and study the dependence of physical quantities on this time. In these cases, CV and CA show a plateau, where complexity is nearly constant, at early times $t_{w}\ll t_{*}$. Next, the holographic conjectures approach a linear growth at late times $t_{w}\gg t_{*}$ The plateau becomes longer when the shock is inserted at earlier times or when the perturbation has smaller energy, showing signatures of scrambling characteristic to chaotic systems. This is another manifestation of the switchback effect. The scrambling time can be extracted by studying the plateau size, and reads $t_{*}\approx\frac{1}{2\pi T}\log(M/T)$, as expected. Applying the CV conjecture, the results are summarized in Fig. 11. Furthermore, one can study the precise early $t_{w}$ dependence of holographic complexity. A careful analysis (summarized in Fig. 27 of [43]) shows that the complexity grows exponentially $\mathcal{C}(t_{w})\approx\exp(2\pi/\beta(t_{w}-t_{*}))$, with a characteristic exponent which is exactly equal to the Lyapunov exponent of maximally chaotic systems (see $\lambda_{\rm OTOC}$ in the key notion 6) [5]. Remarkably, all these behaviors of holographic complexity agree with Eq. (11), obtained from the simple circuit models presented in Sec. 4.1. The matching is evident by comparing Figs. 11 and 5. Finally, let us mention reference [43] extends previous studies to cases where the shocks have energies above the thermal scale.
@@
Figure 11: Volume complexity $\Delta\mathcal{C}_{V}$, compared to the value $\Delta\mathcal{C}_{V,NS}$ in vacuum AdS, as a function of the boundary insertion time $-t_{w}$ of the shock wave. $S_{1}$ is the thermal entropy of the black hole solution before the shock wave insertion; $T_{2}$ the Hawking temperature after the insertion of the shock. Picture taken from reference [43].

### v2

@@ 7 Quantum complexity and space-time: a more concerted approach
Connections between geometry and entanglement were first proposed by Ryu and Takayangi [5] and were made covariant by Hubeny, Rangamani and Takayanagi [237]. These authors showed that the entanglement entropies of subregions of the theory dual to an AdS spacetime are geometrized by the areas of certain extremal surfaces of the dual geometry. These developments led to major insights into the holographic duality, for example elucidating constraints on the entanglement patterns in quantum states that admit gravitational dual descriptions [238]. An analogy between the geometry of AdS spacetimes dual to conformal field theories and the graphical structure of tensor networks for constructing quantum states also provided an impetus, offering a schematic understanding of the hierarchical pattern of short and long range entanglement in holographic states [239]. Such developments motivated an important proposal by van Raamsdonk that spacetime is actually knitted together by the underlying patterns of quantum entanglement [240]. Subsequent developments have illustrated many constraints on quantum states with holographic duals including new inequalities [238] and properties such a quantum error correction [241] (see [242] for a review) that have a bearing on the complexity of constructing such states from a simple reference.
Others proceeded to study the dynamics of thermalization via the AdS/CFT correspondence, especially considering the evolution of entanglement and the corresponding dual geometric quantities after quantum quenches [147, 148, 149, 243, 150, 244]. In all these cases, an initial rise of entanglement entropy ends in saturation after a thermalization time. These studies also led to an examination of the time evolution of entanglement entropy in thermo-field double (TFD) states dual to eternal black holes, i.e., to ERBs between two asymptotically AdS universes [245, 246]. The result, computed in terms of certain extremal surfaces passing through the ERB, shows that entanglement entropy grows with time in a TFD state until it saturates at equilibrium [245]. However, the black hole interior keeps growing long after the entanglement saturates. Therefore, Susskind suggested that the growth of the black hole’s interior should be encoded in the AdS/CFT correspondence by a dual quantum mechanical quantity that increases for a time exponential in the system’s size. What quantity could conceivably keep increasing after a system reaches thermal equilibrium? It is known in computer science that some computations carried out by the action of local operators can require exponentially long time or exponentially many actions of the operators, quantified as the “complexity” of the resulting state. If we think of time evolution as a computation performed by the action of a local Hamiltonian, one might guess that the complexity of the evolving quantum state, quantified in some way, may be related to the size of the black hole’s interior [247].
As we discussed in Secs. 4, 5, and 6, circuit complexity, Nielsen’s complexity, and Krylov/spread complexity all show a long period of linear growth for chaotic systems and also display the switchback effect (see discussion in Sec. 6.3.2). Therefore, any dual geometric object must also show these dynamics. Consistently with this, the initial conjecture was that the volume of the ERB wormhole is dual to some notion of time-evolving complexity of the dual Thermo-field Double state [247]. However, as we discuss in Sections 7.1 and 7.2, a number of different geometric quantities all reproduce linear growth and the switchback effect [25, 26, 27, 28, 29, 30]. We also discuss a proposal in the case of two-dimensional gravity for a non-perturbative definition of the size of the ERB wormhole [31] that saturates at times exponential in the black hole entropy, reproducing the expected late time saturation of complexity (Secs. 4,5 and 6). Strictly speaking, the various candidate holographic duals of complexity described in Sec. 7.1 all have divergences that require regulation. We understand these divergences as arising in the dual CFT from the continuum nature of the theory, just like the divergences of entanglement entropy. We explain these considerations in Sec. 7.3. Proposed extensions of the ideas to subregions, geometries with defects, and black holes in de Sitter space are summarized Sec. 7.4.
All the geometric quantities proposed as duals to some notion of complexity in Sec. 7.1 are diffeomorphism-invariant, and hence we expect them to map onto precise gauge-invariant quantities in the dual theory. Is there a precise map to some particular notion of complexity? In other words, can we go beyond a qualitative agreement with the linear growth and switchback effect seen in any reasonable definition of complexity, to write an equation between gravitational and CFT quantities? Very recently, precise relations have been found between the size of an Einstein-Rosen Bridge in two-dimensional Jackiw-Teitelboim gravity and the spread complexity of a dual double-scaled SYK model, matching both the linear early time classical growth and the expected late time quantum saturation [32, 248, 34, 249, 35]. There are also equalities relating the rate of change of Krylov/spread complexity and the momentum of particles in AdS spacetime. These results are reviewed in Sec. 7.5. Sec. 7.6 discusses some ideas for relating Nielsen’s complexity to dual geometries. Despite these findings, it is difficult to match geometrical observables in the gravitational setting with quantum complexity in field theory. This difficulty led to questions about the foundations of quantum computation in the context of the AdS/CFT correspondence. The existence of pseudorandom states, discussed in Sec. 7.7, can be used to argue that some entries in the holographic dictionary will be hard to compute, or alternatively, that quantum gravity efficiently solves problems that will require exponentially long times for a quantum computer. This unexpected line of research was revealed through the study of quantum complexity.
@@ 7.1 Holographic complexity proposals
We begin by defining the holographic complexity proposals (see Fig. 10).
@@
The complexity equals volume (CV) proposal [25, 6] declares that holographic complexity is proportional to the maximal volume $\mathcal{V}$ of a codimension-one spacelike surface $\mathcal{B}$ anchored at the $(d-1)$-dimensional boundary slice $\Sigma_{\rm CFT}$ where a CFT state is defined (see fig. 10(a)), i.e.,  [^27] In the context of CV conjecture, Ref. [250] proposed to measure this geometric observable in terms of the maximal time from the horizon to the final slice (approached in the late stages of evolution). This notion also shares many similarities with the standard CV proposal, and at the same time removes certain ambiguities between small and large black holes in AdS spacetime.
@@
where $\ell_{\rm bulk}$ is a length scale introduced to maintain dimensional consistency. In the case of AdS spacetime, $\ell_{\rm bulk}$ is often identified with the AdS radius $L$. The maximal surface anchored on both boundaries of an eternal black hole penetrates the black hole interior, but avoids the singularity and regions with high curvature. This makes the maximal volume reliable in the semiclassical approximation, without the need to make further assumptions.
@@
Each of these proposals presents ambiguities. CV and CV2.0 require a length scale $\ell_{\rm bulk}$ in the holographic dictionary. The CA proposal contains contributions from the null boundaries of the WDW patch, whose normal vectors depend on an arbitrary normalization constant. The last dependence can be removed by including certain counterterms in the gravitational action, but those in turn come equipped with their own (arbitrary) length scale, which we denote $L_{\rm ct}$ [251]. When studying black holes, the CA and CV2.0 proposals probe regions near the black hole singularity. A delicate cancellation between a diverging metric factor and the vanishing volume of the sphere in the angular directions near the singularity leads to a result which does not diverge. Nevertheless, the validity of the formula in the semiclassical approximation is uncertain since the region near the singularity is not well understood in this limit.
The above ambiguities were taken as features of the holographic complexity proposal, rather than bugs, since computational complexity is also subject to several ambiguities: for instance, the choice of allowed gates in a circuit, the tolerance to reach a target state, the cost of gates etc., see Sec. 4. This perspective was taken a step further with a recent proposal that went under the name complexity equals anything (CAny), i.e., complexity may be dual to any one out of an infinite class of geometric observables. CAny requires a codimension-zero bulk region $\mathcal{M}$ with future (past) boundary $\Sigma_{+}(\Sigma_{-})$ anchored at the CFT slice $\partial\Sigma_{\pm}=\Sigma_{\rm CFT}$. One defines the following functional (see fig. 10(c)) [29, 30, 47]
@@
where the shape of the boundaries $\Sigma_{\pm}$ is varied. When more than one solution exists, the one with maximal $W_{G_{2},F_{2,\pm}}$ is chosen. Once the unique codimension-zero region (¯,Σ¯±)¯subscript¯$(\bar{\mathcal{M}},\bar{\Sigma}_{\pm})$ solving Eq. (116) is identified, holographic complexity is computed by the observable
$\mathcal{C}_{\rm Any}[G_{1},F_{1,\pm},\bar{\mathcal{M}}](\Sigma_{\rm CFT})=\frac{1}{G_{N}\ell_{\rm bulk}^{2}}\int_{\bar{\mathcal{M}}}d^{d+1}x\,\sqrt{-g}\,G_{1}(g_{\mu\nu})$ (117)
@@
where $G_{1},F_{1,\pm}$ are scalar functions. In summary, the CAny conjecture consists of two steps: an extremization governed by the function (115) to select the relevant bulk region in the geometry, followed by the computation of the gravitational observable in Eq. (117). CAny defines a large class of physical observables, compatible with qualitative features of complexity such as a linear growth and the switchback effect, that can be built out of the building blocks available in the gravitational system.
@@
CV and CV2.0 are recovered by making different terms in Eq. (118) dominant. The CV prescription is recovered by keeping $\alpha_{-},\alpha_{\rm B}$ constant while performing the limit $\mathcal{C}_{V}=\lim_{\alpha_{+}\rightarrow\infty}\left(\alpha_{+}^{-1}\mathcal{C}_{\rm gen}\right)\,.$ In other words, only the future codimension-one surface $\Sigma_{+}$ survives, the extremization problem imposes that the slice has maximal volume, and the physical observable is the volume itself. To recover CV2.0, one fixes the bulk coefficient $\alpha_{\rm B}$ and computes $\mathcal{C}_{2.0V}=\lim_{\alpha_{\pm}\rightarrow 0}\left(\alpha_{B}^{-1}\mathcal{C}_{\rm gen}\right)\,.$ In this limit, the boundary surfaces $\Sigma_{\pm}$ are pushed towards the null surfaces originating from the CFT slice $\Sigma_{\rm CFT}$, thus composing the WDW patch. One can also recover CA from CAny, but now the functionals (115) and (117) should differ from each other. Specifically, one needs to extremize $\mathcal{C}_{\rm gen}$ in Eq. (118) in the limit $\alpha_{\pm}\rightarrow 0$ to identify the WDW patch, but use the gravitational action (113) as the observable which computes the complexity.
@@ 7.2 Time-dependent properties of the holographic proposals
In this subsection, we show that all the holographic complexity proposals defined in Sec. 7.1 universally exhibit the behaviors required for a correspondence with complexity, i.e., linear growth at late times and the switchback effect in a black-hole background.
@@
where $r_{h}$ is the horizon radius. This two-sided geometry is dual to the thermofield-double state $|\mathrm{TFD}\rangle$ of two decoupled boundary CFTs (see Eq. (5)). The isometries of the background (119) allow us to express the complexity observable $\mathcal{C}_{\rm Any}$ in Eq. (117) as follows,
@@
where $\sigma$ is an intrinsic radial parameter, $V_{x}=\int d^{d-1}x$ is the transverse volume, and $v=t+r^{*}(r)$ is an infalling null coordinate defined in terms of $dr^{*}=dr/f(r)$.
Since $v$ is cyclic for $\mathcal{L}_{\varepsilon}$ in Eq. (120), [^28] This means that $\mathcal{L}_{\varepsilon}$ depends on $\dot{v}\coloneqq dv/d\sigma$, but not on $v$. the conjugate momenta $P_{v}^{\varepsilon}=\partial\mathcal{L}_{\varepsilon}/\partial\dot{v}$ are conserved along the profiles of $\Sigma_{\varepsilon}$. The late time regime $t\rightarrow\infty$ is achieved when each of the codimension-one surfaces $\Sigma_{\pm}$ hug a final slice at constant radial coordinate, such that the momenta approach constant values, up to exponentially-suppressed terms. This leads to the linear growth [29, 30]
@@
where $W_{L}$ are low-energy perturbations acting on the left boundary. These perturbations carry energies of the order of the black hole’s temperature (thermal-scale perturbations), which is assumed to be much smaller than the black hole’s mass. Let us assume that $|t_{k+1}-t_{k}|>t_{*}$, where $t_{*}$ is the scrambling time, and the label $k=0,1,\dots,n+1$ in the inequalities includes the boundary times via the definitions $t_{0}=-t_{R}$ and $t_{n+1}=t_{L}$. We also assume that, in the sequence of times $t_{0},\ldots t_{n+1}$ there are $n_{\rm sb}$ switchbacks (also called time folds): in other words, the arguments of the absolute values $|t_{k+1}-t_{k}|$ change sign $n_{\rm sb}\leq n$ times. To keep the bulk solutions as simple as possible, we assume that the perturbations are approximately spherically symmetric. The bulk geometry dual to the perturbed state (122) can be modeled by a black hole with $n\geq 1$ shock waves produced by null matter. In this setting, any CAny observable in the late time limit reads [30]
@@
The contribution $-2n_{\rm sb}t_{*}$ reflects the spreading of the perturbations into the system during a time scale measured by the scrambling time $t_{*}=\frac{1}{2\pi T}\log\left(M/E\right)\approx\frac{1}{2\pi T}\log S$, where $E\ll M$ is the energy of the shock, and $S$ the black hole’s entropy. This characterizes the switchback effect for the CAny observables. At late boundary times $(t_{L},t_{R})$, holographic complexity will grow linearly with the sum of the boundary times $t_{L}+t_{R}$, but with a delay induced by the scrambling time of the perturbation. This phenomenon was already observed in earlier studies of the CV, CV2.0 and CA proposals [25, 27, 28]. This is not a surprise. As we said, those proposals can be obtained from CAny by appropriate limiting procedures.
@@
Let us discuss other refined properties of the holographic conjectures, associated with early and intermediate times compared to the scrambling time and the inverse temperature scale. It was originally speculated that the black hole mass $M$ should provide an upper bound on the complexity rate $d\mathcal{C}/dt\leq 2M/\pi$, based on Lloyd’s bound on the computational speed [256]. This statement was supported by the the full time evolution of CV in an eternal AdS black hole background, since the volume grows at a monotonically increasing rate that approaches a constant from below at late times (e.g., see Fig. 7 in [257]). This constant exactly saturates Lloyd’s bound for large black holes. However, the rate of CA reaches a maximum at a finite time before approaching from above a constant final value saturating Lloyd’s bound (e.g., see Fig. 22 in [257]). Therefore, the intermediate regime in the evolution of CA violates the bound. Other counterexamples have also been found [258, 259, 260, 261, 262, 263, 264, 265, 266, 267]. It was argued in [268] that a black hole should be modeled by the composition of simple gates (i.e., close to the identity), and is thus incompatible with the hypotheses used to derive the Lloyd’s bound. Overall, the adaptation of Lloyd’s bound to holographic complexity is an open problem.
@@
Next, one can study the reaction of the holographic observables to shock waves corresponding to the insertion of null matter at arbitrary boundary times [253, 54]. Earlier, we assumed that the insertion times $t_{k}$ of the perturbations (which included the boundary times $t_{L},t_{R}$) were separated by an interval larger than $t_{*}$ (see the discussion below equation (122)). Here we will not make this assumption. For simplicity, let us fix $t_{L}=t_{R}=0$. We consider the case of a single time fold at the (negative) time $-t_{w}$ when the shock is inserted, and study the dependence of physical quantities on this time. In these cases, CV and CA show a plateau, where complexity is nearly constant, at early times $t_{w}\ll t_{*}$. But these quantities approach linear growth at later times $t_{w}\gg t_{*}$. The plateau becomes longer when the perturbation has smaller energy, showing signatures of scrambling in chaotic systems. [^29] Alternatively, one can study the full dependence of CV and CA on the boundary times, e.g., see Fig. 11 and Fig. 17 in Ref. [54]. In this case, holographic complexity admits a plateau region around $t_{L}=t_{R}=0$, which becomes longer when the shock is inserted at earlier times. This is another manifestation of the switchback effect. The scrambling time can be extracted by studying the plateau size, and turns out to be $t_{*}\approx\frac{1}{2\pi T}\log(M/T)$ (where $T$ is the Hawking temperature), as expected. Applying the CV conjecture, the results are summarized in Fig. 11. Furthermore, one can study the precise early $t_{w}$ dependence of holographic complexity. A careful analysis (summarized in Fig. 27 of [54]) shows that the complexity grows exponentially $\mathcal{C}(t_{w})\approx\exp(2\pi(t_{w}-t_{*})/\beta)$, with a characteristic exponent which is exactly equal to the Lyapunov exponent of maximally chaotic systems (see $\lambda_{\rm OTOC}$ in the key notion 6) [10]. Remarkably, all these behaviors of the quantities proposed as duals to holographic complexity agree with Eq. (11), obtained from the simple circuit models presented in Sec. 4.1. The matching is evident by comparing Figs. 11 and 5. Finally, [54] extends previous studies to cases where the shocks have energies above the thermal scale.
@@
Figure 11: Volume complexity $\Delta\mathcal{C}_{V}$ in three bulk (two boundary) dimensions, compared to the value $\Delta\mathcal{C}_{V,NS}$ of the complexity in the Neveu-Schwarz vacuum, as a function of the boundary insertion time $-t_{w}$ of the shock wave. $S_{1}$ is the thermal entropy of the black hole solution before the shock wave insertion; $T_{2}$ the Hawking temperature after the insertion of the shock. Picture taken from Fig. 19 of reference [54]. From bottom (green) to top (orange), the colors correspond to increasing value of the energy of the shockwave as follows: $E_{\text{shock}}/(2M_{1})=1.5\text{ (orange)},0.625\text{ (pink)},0.105\text{ (blue)},10^{-4}\text{ (red)},10^{-8}$ (green).